OpenStudy (anonymous):

The formula to find the perimeter of a rectangle is: P = 2W + 2L. To solve for " W " in this formula, which steps should be followed? A.) Subtract 2L from both sides of the equation, then subtract 2 from both sides of the equation. B.) Subtract 2L from both sides of the equation, then divide both sides of the equation by 2. C.) Divide both sides of the equation by 2L, then subtract 2 from both sides of the equation. D.) Divide both sides of the equation by 2L + 2

OpenStudy (anonymous):

i think is b

OpenStudy (dumbcow):

uggh i hate these type of questions .... why dont they just say..solve for "W" then if you get it right ,obviously you know the steps

OpenStudy (anonymous):

p-2l=2w+2l-2l p-2l=2w p-2l/2=2w/2 p-2l/2=w

OpenStudy (dumbcow):

yeah its B